# Know What You Don’t Know: Uncertainty Calibration of Process Reward Models

### v1

@@ Abstract
Process reward models (PRMs) play a central role in guiding inference-time scaling algorithms for large language models (LLMs). However, we observe that even state-of-the-art PRMs can be poorly calibrated and often overestimate success probabilities. To address this, we present a calibration approach—performed via quantile regression—that adjusts PRM outputs to better align with true success probabilities. Leveraging these calibrated success estimates and their associated confidence bounds, we introduce an instance-adaptive scaling (IAS) framework that dynamically adjusts the inference budget based on the estimated likelihood that a partial reasoning trajectory will yield a correct final answer. Unlike conventional methods that allocate a fixed number of reasoning trajectories per query, this approach successfully adapts to each instance and reasoning step when using our calibrated PRMs. Experiments on mathematical reasoning benchmarks show that (i) our PRM calibration method successfully achieves small calibration error, outperforming the baseline methods, (ii) calibration is crucial for enabling effective adaptive scaling, and (iii) the proposed IAS strategy reduces inference costs while maintaining final answer accuracy, utilizing less compute on more confident problems as desired. Our implementation is available at https://github.com/azizanlab/instance-adaptive-scaling.
@@ 1 Introduction
Miscalibration and overconfidence of PRMs can limit their utility beyond simply ranking partial reasoning steps. We highlight three important applications where calibrated PRMs are desirable: (1) providing interpretable uncertainty estimates to monitor LLM outputs during generation [53], (2) identifying when to say “I don’t know” [see e.g., 43, 6] or backtrack based on a low predicted probability of success, and (3) assessing the current problem difficulty or likelihood of success to adaptively adjust the inference-time compute budget. These applications motivate a fundamental question: Can we calibrate off-the-shelf PRMs and leverage them to enhance inference-time scaling methods?
@@
Figure 1: Why should problem-solving strategies adapt to task difficulty? Simple problems can often be solved through quick, intuitive solutions, whereas harder ones require extended, deliberate reasoning steps. Accordingly, it is desirable for LLMs to adjust their compute usage adaptively. Fixed-budget methods like best-of-$N$ are inefficient in this context: they either waste computation on easy tasks or fail to allocate sufficient resources to solve harder ones accurately.
@@
In this paper, we introduce a pipeline for improving the calibration of any off-the-shelf PRMs, allowing their scores to more accurately reflect the uncertainty that a given LLM will reach the correct answer. We begin by showing that standard calibration techniques, such as temperature scaling [15], are inadequate for calibrating PRMs. To address this, we introduce a quantile regression [21] based scheme that reduces the calibration error of PRM scores. The resulting model predicts success probabilities, together with confidence bounds, for every query and intermediate reasoning step. Our method collects intermediate reasoning steps generated by any LLM and completes them by independently rolling out full responses. This Monte Carlo rollout provides an empirical estimate of the true success probability for each prefix trajectory. Given this data, we fine-tune the PRM by replacing its prediction head with a quantile regression model, optimizing only this layer for efficiency. The resulting model produces reliable success probability estimates when evaluating intermediate steps in multi-step reasoning tasks.
@@
We propose to use these now-calibrated PRMs to enable instance-adaptive scaling (IAS), a framework where test-time compute is adaptively scaled based on the calibrated PRM’s assessment of difficulty/likelihood of success. Beyond saving compute, this concept has the potential to reduce overall latency for the user waiting for a response. The intuition mirrors how humans solve problems: we spend more time on difficult questions and allocate more effort to promising solution paths (see Figure 1). Similarly, IAS guides LLMs to invest more computation in challenging or high-potential reasoning paths. We focus on two widely used inference-time scaling methods: best-of-$N$ (BoN) and beam search (BS). We prove that, for any given query and reasoning trajectory, the calibrated reward score can estimate the minimum number of additional trajectories required to produce at least one correct answer. In short, our key contributions are:
@@
  • We address the overestimation of success probabilities in existing process reward models by introducing a calibration approach using quantile regression.
@@
  • We propose an instance-adaptive inference-time scaling strategy that dynamically allocates computation by leveraging calibrated PRMs.
@@ 1.1 Related Work
Process reward models Process reward models are specialized tools used in inference-time scaling that provide a step-by-step verification of LLMs’ reasoning process, rather than evaluating only the final outcome [57]. Typically, PRMs are trained on datasets labeled at each reasoning step, either by humans or automated approaches like Monte Carlo rollouts and LLM-based judges. For example, Qwen-PRM [51] utilizes consensus-filtered labels from combined human and automated sources, achieving state-of-the-art accuracy [41], while Shepherd-PRM [47] effectively relies on purely automated labeling despite lower precision in detecting errors.
@@
2 Preliminaries
2.1 Model Calibration
Calibration error metrics assess whether a model’s predicted probabilities are consistent with observed accuracy. Two widely used calibration metrics are the Brier score [2] and the expected calibration error (ECE) [30]. Brier score quantifies the mean-squared difference between predicted probabilities $\hat{y}$ and ground truth labels $y$:
$\text{Brier}\;=\;\frac{1}{|\mathcal{D}|}\,\sum_{l=1}^{|\mathcal{D}|}\bigl{(}\hat{y}^{(l)}-y^{(l)}\bigr{)}^{2}.$
$\text{ECE}\;=\;\sum_{b=1}^{B}\frac{|\mathcal{I}_{b}|}{|\mathcal{D}|}\,\bigl{|}\operatorname{conf}(\mathcal{B}_{b})-\operatorname{acc}(\mathcal{B}_{b})\bigr{|},$
where $\operatorname{conf}(\mathcal{B}_{b})=\tfrac{1}{|\mathcal{I}_{b}|}\sum_{l\in\mathcal{I}_{b}}\hat{y}^{(l)}$ is the average confidence in bin $b$ and $\operatorname{acc}(\mathcal{B}_{b})=\tfrac{1}{|\mathcal{I}_{b}|}\sum_{l\in\mathcal{I}_{b}}y^{(l)}$ is the empirical accuracy. Variants of ECE include average calibration error (ACE) [31], which uses equal weighting across bins.
These different metrics all have advantages and disadvantages, e.g. Brier score does not (usually) go to zero with perfect calibration (while ECE does), and the binning-based metrics can perhaps unfairly reward predictor collapse—these go to zero on a dataset if the predictor simply returns the same dataset average performance regardless of input; since all predictions would fall in the same bin.
2.2 Inference-time Scaling with Process Reward Models
We recall how reward models are used by standard inference-time scaling methods. Let $\mathsf{LLM}$ denote the language model that generates responses, and let $\mathsf{PRM}$ denote the process reward model used to evaluate the quality of the outputs produced by the $\mathsf{LLM}$. Given a query $q$, the $\mathsf{LLM}$ generates a multi-step reasoning trajectory $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})$, where $x_{i}$ represents the $i$-th reasoning step and $T$ is the total length of the trajectory. We denote the prefix of the trajectory up to step $t$ as $\mathbf{x}_{1:t}$. Below, we recap two well-established and foundational PRM-based inference-time scaling methods.
Best-of-$N$ (BoN) [3]: Given a query $q$, we first use the $\mathsf{LLM}$ to generate $N$ complete trajectories
@@ 2.2 Inference-time Scaling with Process Reward Models
Then we apply the $\mathsf{PRM}$ to assign a score to each trajectory: $r^{(i)}=\mathsf{PRM}\left(q,\mathbf{x}^{(i)}\right)$. The final output is the trajectory with the highest reward $\mathbf{x}^{(i^{*})}$ where $i^{*}=\arg\max_{i}r^{(i)}$.
Beam Search (BS) [40]: BS iteratively builds reasoning trajectories by alternating between generation and evaluation, rather than generating full sequences in a single pass. At step $t$, each of the $K$ surviving partial trajectories $\mathbf{x}_{1:t}^{(i)}$ is extended by one reasoning step. For each trajectory, the $\mathsf{LLM}$ generates $M$ candidate next steps: $\mathbf{x}_{t+1}^{(i,j)}\sim\mathsf{LLM}\left(q,\mathbf{x}_{1:t}^{(i)}\right)$ where $j=1,\ldots,M$ denotes different continuations for the $i$-th trajectory. Then the $\mathsf{PRM}$ evaluates each of the $N=K\times M$ extended trajectories: $r_{t}^{(i,j)}=\mathsf{PRM}\left(q,\mathbf{x}_{1:t+1}^{(i,j)}\right)$. Finally, among all $N$ candidates, only the top $K$ partial trajectories with the highest scores are kept for the next step. This pruning step ensures the search remains computationally tractable while focusing on promising reasoning paths.
3 Calibration and Process Rewards
PRMs are typically trained to quantify the quality of (partial) model outputs, i.e., their correctness, how well they advance the reasoning towards the solution to the query, and/or how well they align with human preferences. We highlight an additional, underexplored perspective: when normalized, PRM scores can be interpreted as an estimate of the probability that continuing from the current output will ultimately lead to a desirable outcome, such as producing a correct final answer in a multi-step reasoning task [47]. We observe that state-of-the-art PRMs often overestimate the chance of eventual success when interpreted in this way. We define success probability as follows:
Definition 1.
Given a query $q$ and a partial trajectory $\mathbf{x}_{1:t}$ generated so far, we define its success probability under a language model $\mathsf{LLM}$ as the probability that continuing this trajectory will ultimately lead to a correct final answer:
$p\;\triangleq\;\Pr\bigl{(}\mathsf{LLM}\text{ eventually yields a correct answer}\,\bigm{|}\,q,\mathbf{x}_{1:t}\bigr{)}.$
Since we have partial outputs, there is inherent uncertainty of future success (i.e. $p\in(0,1)$), unlike single-turn settings where an output can definitively be evaluated for correctness. Note that this is a significant difference between our work and typical prior work on calibrated certainty estimates for LLMs, and is crucial to our quantile-based approach being feasible. That said, calibration metrics like Brier score and ECE still apply.
Why calibrate PRMs? While we will show that a well-calibrated PRM is desirable for several important tasks, calibration is not required for the canonical task of ranking for BoN or BS, since relative ordering is all that matters. Specifically, any monotonic transformation of the true success probability preserves the ranking but changes calibration.
As noted in the introduction, however, calibrated probabilities are essential for interpretability, systematic safety monitoring, and efficient budget allocation. In particular, calibration enables instance-adaptive scaling (IAS), where we dynamically adjust the number of trajectories to keep based on the estimated likelihood of success. In this setting, accurate probability estimates are essential for balancing efficiency and performance: a well-calibrated PRM allows us to reduce computations and maintain output quality. Furthermore, it is important here to use a PRM, rather than a traditional reward model, as we are evaluating either partial (BS) or non-existent (BoN) generations when choosing the budget.
@@ 3 Calibration and Process Rewards
Figure 2: Histogram of signed deviation (i.e., estimation error) for Qwen-PRM and Shepherd-PRM evaluated on the MATH500 (in-distribution) and AIME24-25 (out-of-distribution) datasets. Positive error indicates overestimation. As shown, PRMs tend to be optimistic (e.g., exhibiting higher density on the right side compared to the left and/or forming a noticeable peak around $1.0$). This becomes particularly pronounced for weaker models and more challenging, out-of-distribution problems.
3.1 Instance-Adaptive Inference-Time Scaling

### v2

@@ Abstract
Process reward models (PRMs) play a central role in guiding inference-time scaling algorithms for large language models (LLMs). However, we observe that even state-of-the-art PRMs can be poorly calibrated. Specifically, they tend to overestimate the success probability that a partial reasoning step will lead to a correct final answer, particularly when smaller LLMs are used to complete the reasoning trajectory. To address this, we present a calibration approach—performed via quantile regression—that adjusts PRM outputs to better align with true success probabilities. Leveraging these calibrated success estimates and their associated confidence bounds, we introduce an instance-adaptive scaling (IAS) framework that dynamically adjusts the compute budget based on the estimated likelihood that a partial reasoning trajectory will yield a correct final answer. Unlike conventional methods that allocate a fixed number of reasoning trajectories per query, this approach adapts to each instance and reasoning step when using our calibrated PRMs. Experiments on mathematical reasoning benchmarks show that (i) our PRM calibration method achieves small calibration error, outperforming the baseline methods, (ii) calibration is crucial for enabling effective IAS, and (iii) the proposed IAS strategy reduces inference costs while maintaining final answer accuracy, utilizing less compute on more confident problems as desired.
@@ 1 Introduction
Miscalibration and overconfidence of PRMs can limit their utility beyond simply ranking partial reasoning steps. We highlight three important applications where calibrated PRMs are desirable: (1) providing interpretable uncertainty estimates to monitor LLM outputs during generation [58], (2) identifying when to say “I don’t know” [see, e.g., 46, 7] or backtrack based on a low predicted probability of success, and (3) assessing the current problem difficulty or likelihood of success to adaptively adjust the inference-time compute budget. These applications motivate a fundamental question: Can we calibrate off-the-shelf PRMs and leverage them to enhance inference-time scaling methods?
@@
Figure 1: Why should problem-solving strategies adapt to task difficulty? Simple problems can often be solved through quick, intuitive solutions, whereas harder ones require extended, deliberate reasoning steps. Accordingly, it is desirable for LLMs to adjust their compute usage adaptively. Fixed-budget methods like best-of-NN are inefficient in this context: they either waste computation on easy tasks or fail to allocate sufficient resources to solve harder ones accurately.
@@
In this paper, we introduce a pipeline for improving the calibration of any off-the-shelf PRMs, allowing their scores to more accurately reflect the uncertainty that a given LLM will reach the correct answer. We begin by showing that standard calibration techniques, such as temperature scaling [16], are inadequate for calibrating PRMs. To address this, we introduce a quantile regression [22] based scheme that reduces the calibration error of PRM scores. The resulting model predicts success probabilities, together with confidence bounds, for every query and intermediate reasoning step. Our method collects intermediate reasoning steps generated by a given LLM and completes them by independently rolling out full responses. This Monte Carlo rollout provides an empirical estimate of the true success probability for each prefix trajectory. Given this data, we fine-tune the PRM after replacing its prediction head with a quantile regression model. The resulting model produces reliable success probability estimates when evaluating intermediate steps in multi-step reasoning tasks.
@@
We use these now-calibrated PRMs to enable instance-adaptive scaling (IAS), a framework where test-time compute is adaptively scaled based on the calibrated PRM’s assessment of difficulty/likelihood of success. Beyond saving compute, this concept has the potential to reduce overall latency for the user waiting for a response. The intuition mirrors how humans solve problems: we spend more time on difficult questions and allocate more effort to promising solution paths (see Figure 1). Similarly, IAS guides LLMs to invest more computation in challenging or high-potential reasoning paths. We focus on two widely used inference-time scaling methods: best-of-$N$ (BoN) and beam search (BS). We prove that, for any given query and reasoning trajectory, the calibrated reward score can estimate the minimum number of additional trajectories required to produce at least one correct answer. In short, our key contributions are:
@@
  • We address the overestimation of success probabilities in existing PRMs by introducing a calibration approach using quantile regression.
@@
  • We propose an instance-adaptive inference-time scaling strategy that dynamically allocates compute budgets by leveraging calibrated PRMs.
@@ 1.1 Related Work
Process reward models. Process reward models are specialized tools used in inference-time scaling that provide a step-by-step verification of LLMs’ reasoning process, rather than evaluating only the final outcome [62]. Typically, PRMs are trained on datasets labeled at each reasoning step, either by humans or automated approaches like Monte Carlo rollouts and LLM-based judges. For example, Qwen-PRM [55] utilizes consensus-filtered labels from combined human and automated sources, achieving state-of-the-art accuracy [44], while Shepherd-PRM [50] effectively relies on purely automated labeling despite lower precision in detecting errors.
@@
2 Preliminaries and Notation: Inference-Time Scaling with PRMs
We recall how reward models are used by standard inference-time scaling methods. Let $\mathsf{LLM}$ denote a language model that generates responses, and let $\mathsf{PRM}$ denote a process reward model used to evaluate the quality of the outputs produced by $\mathsf{LLM}$. Given a query $q$, $\mathsf{LLM}$ generates a multi-step reasoning trajectory $\mathbf{x}=(x_{1},x_{2},\ldots,x_{T})$, where $x_{i}$ represents the $i$-th reasoning step and $T$ is the total length of the trajectory. We denote the prefix of the trajectory up to step $t$ as $\mathbf{x}_{0:t}$ (in particular, $x_{0}:=$“ ” and $x_{0:0}$ is an empty reasoning sequence). Below, we recap two standard PRM-based inference-time scaling methods.
Best-of-NN (BoN) [4]: Given a query $q$, we first use $\mathsf{LLM}$ to generate $N$ complete trajectories
@@ 2 Preliminaries and Notation: Inference-Time Scaling with PRMs
Then we apply $\mathsf{PRM}$ to assign a score to each trajectory: $r^{(i)}=\mathsf{PRM}\left(q,\mathbf{x}^{(i)}\right)$. The final output is the trajectory with the highest reward $\mathbf{x}^{(i^{*})}$ where $i^{*}=\arg\max_{i}r^{(i)}$.
Beam Search (BS) [43]: BS iteratively builds reasoning trajectories by alternating between generation and evaluation, rather than generating full sequences in a single pass. At step $t$, each of the $K$ surviving partial trajectories $\mathbf{x}_{0:t-1}^{(i)}$ is extended by one reasoning step. Specifically, for each trajectory, $\mathsf{LLM}$ generates $M$ candidate next steps: $\mathbf{x}_{t}^{(i,j)}\sim\mathsf{LLM}\left(q,\mathbf{x}_{0:t-1}^{(i)}\right)$ where $j=1,\ldots,M$ denotes different continuations for the $i$-th trajectory. Then $\mathsf{PRM}$ evaluates each of the $N=K\times M$ extended trajectories: $r_{t}^{(i,j)}=\mathsf{PRM}\left(q,\mathbf{x}_{0:t}^{(i,j)}\right)$. Finally, among all $N$ candidates, only the top $K$ partial trajectories with the highest scores are kept for the next step. This pruning step ensures the search remains computationally tractable while focusing on promising reasoning paths.
@@
Figure 2: Histogram of signed deviations between PRM rewards (i.e., estimated success probabilities) and ground-truth success probabilities. Ground truth is estimated via Monte Carlo sampling: for each question and partial reasoning step prefix, we use a given LLM to generate multiple completions and compute the empirical success rate. We evaluate Qwen-PRM-7B and Shepherd-PRM-7B on the MATH500 (in-distribution) and AIME24-25 (out-of-distribution) datasets. Positive deviations indicate overestimation. PRMs consistently overestimate success probabilities, as evidenced by the distribution skewing right and/or peaking near $1.0$. This miscalibration is particularly pronounced for weaker completion models and more challenging, out-of-distribution problems.
3 Calibration and Process Rewards
PRMs are typically trained to quantify the quality of (partial) model outputs, i.e., their correctness, how well they advance the reasoning towards the solution to the query, and/or how well they align with human preferences. We highlight an additional, underexplored perspective: when normalized, PRM scores can be interpreted as an estimate of the probability that continuing from the current output will ultimately lead to a desirable outcome, such as producing a correct final answer in a multi-step reasoning task [50]. We observe that state-of-the-art PRMs often overestimate the chance of eventual success when interpreted in this way. We define the success probability as follows.
Definition 1.
Given a query $q$ and a partial trajectory $\mathbf{x}_{0:t}$ generated so far, the success probability $p$ is the probability that autoregressively continuing this trajectory with a stochastic language model (i.e., non-zero temperature), $\mathsf{LLM}$, will ultimately yield a correct final answer.
$p\;$ $\triangleq\;\Pr\bigl(x_{t+1:T}\text{ generated by }\mathsf{LLM}\text{ yields a correct answer}\,\bigm|\,q,\mathbf{x}_{0:t}\bigr)$
Here, we denote the index of a reasoning step by $t$ and the partial reasoning trajectory $\mathbf{x}_{0:t}$ is the sequence of outputs generated by the model from step $1$ to step $t\leq T$.
For the base case at $t=0$, the partial trajectory $\mathbf{x}_{0:0}$ is an empty sequence, representing the state before any reasoning steps have been generated, where only the query $q$ is given. In this case, $p$ essentially estimates the difficulty of the query for $\mathsf{LLM}$.
Since we have partial outputs, there is inherent uncertainty of future success (i.e., $p\in(0,1)$), unlike single-turn settings where an output can definitively be evaluated for correctness. Note that this is a significant difference between our work and typical prior work on calibrated certainty estimates for LLMs, and is crucial to our quantile-based approach being feasible. That said, calibration metrics like Brier score and expected calibration error (ECE) still apply.
The inherent calibration challenge of off-the-shelf PRMs. Off-the-shelf PRMs lack a guarantee of calibration because their training is inherently dependent on a specific policy model. They learn a reward function from reasoning paths (rollouts) generated by one particular model (e.g., Qwen-Math-Instruct), tying their accuracy to that model’s unique generative process and capabilities.
The problem originates from the autoregressive nature of language models, where sequence generation is policy-dependent. The probability of future tokens, $\pi_{\theta}(x_{t+1:T}|x_{0:t})$, is conditioned on the model’s parameters, $\theta$. Consequently, a PRM is only calibrated to the statistical patterns of the policy it was trained with. Using it with a different policy introduces a distributional mismatch, breaking any calibration guarantee.
For instance, a PRM trained on a highly capable 72B model will overestimate the success probability when paired with a weaker 1B model. The PRM is calibrated to the stronger model’s performance and cannot account for the weaker model’s higher tendency to make errors on complex logical steps. This mismatch results in inflated and unreliable confidence scores, as presented in Figure 2.
Why calibrate PRMs? While we will show that a well-calibrated PRM is desirable for several important tasks, calibration is not required for the canonical task of ranking for BoN or BS, since relative ordering is all that matters. Specifically, any monotonic transformation of the true success probability preserves the ranking but changes calibration. As noted in the introduction, however, calibrated probabilities are essential for interpretability, systematic safety monitoring, and efficient budget allocation. In particular, calibration enables instance‑adaptive scaling (IAS; discussed further in Section 4), where we dynamically adjust the number of trajectories to keep based on the estimated likelihood of success. In this setting, accurate probability estimates are essential for balancing efficiency and performance: a well-calibrated PRM allows us to reduce computations and maintain output quality. Furthermore, using a PRM rather than an outcome reward model is critical here, as we need to estimate success probabilities before complete solutions exist. For beam search, we evaluate each partial reasoning prefix to guide search decisions. For best-of-$N$, we evaluate only the question itself (with no prefix) to determine the optimal sampling budget.
3.1 Calibration Data Collection Methodology
Now, we present a calibration-via-finetuning approach to mitigate the aforementioned issues. To calibrate a PRM for a given LLM, we construct a validation dataset through a systematic multi-stage generation and evaluation process. Please see details in Appendix C.1.
Our validation set is defined as $\cup_{q\in\mathcal{Q}_{\mathrm{val}}}\{\mathbf{x}^{(q,1)},\mathbf{x}^{(q,2)},\dots,\mathbf{x}^{(q,N_{\mathrm{val}})}\}$, where $\mathcal{Q}_{\mathrm{val}}$ represents a curated set of validation questions; in our experiments, we sample 500 random questions from MATH training split. For notational simplicity, we omit the superscript $q$ in the subsequent discussion, with the understanding that all operations are performed independently for each question in $\mathcal{Q}_{\mathrm{val}}$.
Stage 1: Initial Trajectory Generation. For each validation question $q\in\mathcal{Q}_{\mathrm{val}}$, we generate $N_{\mathrm{val}}=8$ independent reasoning trajectories using the target LLM. This yields a set of complete reasoning chains: $\{\mathbf{x}^{(1)},\mathbf{x}^{(2)},\dots,\mathbf{x}^{(N_{\mathrm{val}})}\}$. These trajectories represent diverse solution paths the model naturally explores for the given question.
Stage 2: Prefix Extraction and Monte Carlo Rollouts. For each reasoning trajectory $\mathbf{x}^{(i)}=x_{0:T^{(i)}}^{(i)}$ generated in Stage 1, we consider all possible prefix trajectories $x_{0:t}^{(i)}$, where $t\in\{0,1,\dots,T^{(i)}\}$ ranges from the initial step to the final step of the trajectory. For each prefix $x_{0:t}^{(i)}$, we perform a Monte Carlo estimation by generating $N_{\mathrm{MC}}=8$ additional follow-up trajectories. These follow-up trajectories are conditioned on both the original question $q$ and the specific prefix $x_{0:t}^{(i)}$, allowing us to estimate the probability of eventual success from that intermediate reasoning state.
Stage 3: Success Probability Estimation. For each prefix trajectory $x_{0:t}^{(i)}$, we evaluate all $N_{\mathrm{MC}}$ follow-up trajectories to determine correctness. Let $Z^{(i,t)}$ denote the count of correct completions among these follow-ups. We then compute the empirical success probability: $\tilde{p}^{(i,t)}=Z^{(i,t)}/N_{\mathrm{MC}}$, which serves as our ground-truth label for calibrating the PRM’s predicted rewards.
Our calibration datasets, consisting of $\bigl(q,x_{0:t}^{(q,i)},\tilde{p}^{(q,i,t)}\bigr)$ triplets (question, prefix, and empirical success probability), are publicly available for various reasoning LLMs. [^1] https://huggingface.co/datasets/young-j-park/prm_calibration
3.2 Uncertainty-Aware Calibration through Quantile Regression
We seek to fine-tune PRMs to align their predictions with success rates $p^{(i,t)}$ given the input $q$ and generated steps so far $x_{0:t}^{(i)}$. However, predicting success probabilities is inherently uncertain: calibration data is non-exhaustive and model capacity is finite. As with any machine learning model trained via empirical risk minimization (ERM), PRMs make loss-minimizing predictions that approximate the conditional mean of the success probability. [^2] Exactly the conditional mean for MSE loss and approximately so for cross-entropy loss when the posterior variance is low.
While the conditional mean is a reasonable point estimate, it presents a critical problem: for any given test case, there is roughly a 50% chance the true success probability falls below this estimate. This means standard PRM architectures will systematically overestimate success probability for half of all instances. For applications requiring conservative estimation—such as instance-adaptive scaling (IAS), where we allocate computational budget based on estimated difficulty—overestimation is particularly problematic. If we overestimate $p^{(i,t)}$, we allocate insufficient samples, leading to high failure rates on problems that genuinely require more computation.
Instead of predicting the conditional mean, we propose predicting a lower quantile of the posterior distribution over $p^{(i,t)}$. By targeting, say, the 10th percentile, we ensure that for at least 90% of test cases, our estimate is below or equal to the true success probability. This provides a conservative lower bound that protects against underallocation of compute while still enabling meaningful instance-adaptive decisions.
To achieve this, we modify the PRM architecture to perform quantile regression. Specifically, we expand the output dimension of the prediction head to generate multiple predictions—one for each target quantile level $\beta_{n}$ (e.g., 10%, 50%, and 90%)—rather than a single reward value. Details of this architectural modification and initialization are in Appendix C.5.
We then fine-tune the model using a weighted quantile loss (wQL) [23]:
[…]
where $N_{q}$ is the number of quantiles, $\hat{r}^{(\beta_{n})}$ is the modified PRM’s prediction for the $\beta_{n}$-quantile, and $\tilde{p}$ is the empirical success rate. This asymmetric loss penalizes overestimation more heavily for lower quantiles and underestimation more heavily for higher quantiles, encouraging the model to learn the distinct quantile levels. Importantly, training requires only point estimates $\tilde{p}$ from Monte Carlo sampling—we do not need to estimate the full distribution of $p$.
4 Instance-Adaptive Inference-Time Scaling